### v1

@@ 2.3 General Jacobi case
The methods considered for the general Gauss-Jacobi case combine the iterative method of [13] with the asymptotic methods of [11]. The iterative method provides a fast method of computation for a wide range of parameters; however, the accuracy of the weights worsens gradually as the degree increases and the asymptotic methods can be used as a replacement for moderate values of the parameters $\alpha$ and $\beta$.
@@ 2.3.1 Iterative method
The previous theorem gives a procedure to compute successive zeros in the direction of decreasing values of $\Omega(z)$. For a simple application of the method it is convenient that the monotonicity properties of $\Omega(z)$ are available and that the variation of the coefficient $\Omega(z)$ is as small as possible, because the method is exact if $\Omega(z)$ is constant. Different changes of variable can be considered such that the transformed differential equation becomes suitable in this sense. This has to be combined with an adequate method for computing the Jacobi polynomial and its derivative (or related functions [^2] Notice that the fixed point method applies to ODEs in normal form, and that transforming the ODE to this form may require a change of function, as for example in Eq. (1)); in our algorithms, as in [13], the main method considered for this purpose is Taylor series, with the successive derivatives in the series computed by differentiation of the ODE satisfied by Jacobi polynomials (or related functions).
The algorithm for the general Jacobi case combines the use of three different changes of variable. For computing the Jacobi polynomial and its derivative (or related functions) with Taylor series the original variable $x$ is used because the differential equation has polynomial coefficents and it is therefore easy to differentiate. The principal change of variable considered for the applying the fixed point method is $z=\tanh^{-1}(x)$; the coefficient of the Liouville-transformed ODE in normal form in the variable $z$ has even simpler monotonicity properties than that of Eq. (2). Finally, for nodes close to the end points $\pm 1$ the angular variable $\theta=\arccos x$ is used to refine the results obtained with the fixed point method associated to the principal change of variable.
For the particular case of Gauss-Gegenbauer quadrature we have implemented an arbitrary precision Maple algorithm in addition to the Matlab algorithm. The computation of the Gegenbauer polynomial and its derivative (or related functions) by Taylor series is particularly simple due to the symmetry of the problem. The angular variable is not used in this Maple algorithm.
For further details on the iterative method for Gauss-Jacobi quadrature, we refer to [13].
@@ 2.3.2 Asymptotic method
Asymptotic methods have been employed for the computation of Gauss-Jacobi quadratures both in combination with iterative methods as in [14] or as independent iteration-free methods for Gauss-Legendre [1] and Gauss-Jacobi [11]. The Gauss-Hermite and Gauss-Laguerre cases were also described in [9]. In [19], a Riemann-Hilbert analysis is considered for the three classical cases and some extensions.
In our algorithms, the asymptotic computation of Gauss-Jacobi quadratures is based on the two types of asymptotic expansions described in [11]: an expansion in terms of elementary functions and a Bessel-type expansion [4]. The elementary expansion is valid in any fixed compact subinterval of (1,1)(-1,1) (therefore the nodes closest to $\pm 1$ are excluded). Contrarily, the Bessel expansion is uniformly valid in $1+\delta,1]$, where $\delta$ is an arbitrarily small positive number; because we can use the symmetry property $P_{n}^{(\alpha,\beta)}(x)=(-1)^{n}P_{n}^{(\beta,\alpha)}(-x)$ this expansion can be used for computing all the zeros of the Jacobi polynomials in (1,1)(-1,1).
The Bessel-type expansion is the main method of computation in our algorithms. However, the expansion in terms of elementary functions is convenient for computing the nodes with small absolute value in order to maintain relative accuracy, and not only absolute accuracy, similarly as discused ealier for Gauss-Legendre quadrature. In our algorithm, $10\%$ of the nodes are computed with the elementary expansion and the rest with the Bessel-type expansion.
@@ 2.3.3 Combining the iterative and asymptotic methods
Our algorithm computes Gauss-Jacobi quadratures using the iterative method and the asymptotic expansions in complementary regions. The iterative method is fast and globally convergent and can be used for practically any values of the parameters, but with some error degradation as the degree increases. For that reason, it is convenient to replace the iterative method by asymptotic computation for sufficiently large degrees.
@@
The accuracy comparison between the asymptotic and iterative methods is shown in Fig. 2, where accuracy comparisons are performed for $10^{4}$ points randomly and uniformly generated in the intervals $n\in[1,3000]$, $\alpha,\beta\in(-1,15]$. The iterative method, as we will se later, turns out to be around a factor $2$ faster that the asymptotic method and, for this reason, when the accuracy of both methods is similar we prefer to use the iterative method. With this consideration, and given the accuracy checks shown in Fig. 2, a good compromise consist in using the asymptotic method when $n>250$ and $\alpha^{2}+\beta^{2}<4(n-195)/55$, and the iterative method otherwise. To this, we add the restriction $\alpha^{2}+\beta^{2}<2400$ in order to avoid overflow/underflow problems in the computation of the asymptotic expansions (not shown in Fig. 2).
Figure 2: Comparison of the accuracy of the iterative and the asymptotic method for Gauss-Jacobi quadratures for three different measures of accuracy: average relative error (left), maximum relative error (center) and relative total error (right). Denoting by $R$ the ratio of the accuracy of the asymptotic method over the accuracy of the iterative method, we have $R<1$ at the red points. $1\leq R<10$ at the green points and $R\geq 10$ at the blue points. The separation line $\alpha^{2}+\beta^{2}=\min\{\frac{\displaystyle{4}}{\displaystyle{55}}(n-195)\}$ is also shown.
@@ 2.3.4 Testing the combined algorithm
Our algorithm is faster and can be used reliably in a larger region of parameters compared to previous methods, particularly for large values of $\alpha$ and $\beta$. Next we analyze the performance of the combined algorithm in terms of its range of computation, accuracy and speed.
@@
As shown in Fig. 3, our algorithm is able to compute Gauss-Jacobi quadratures parameters for $\alpha,\,\beta\in(-1,1000]$ and $n\leq 10^{6}$, while for larger values of $\alpha$ and $\beta$ the algorithm may fail to produce results. For generating the points for the figure we have considered values of $n$ with $n=\lfloor 10^{6R}\rfloor$ with $R$ uniform pseudorandom numbers in $[0,1]$; $\alpha$ and $\beta$ are also generated by exponentiation of pseudorandom numbers.
@@
The maximum relative error when the iterative method is applied usually takes place at the extremes of the interval, which corresponds with the less significant weights when $\alpha,\beta>-1/2$ (see Eq. (4)). This is expected because these are the last computed weights in the iterative algorithm and they are the closest to the singularities of the ODE (which affects the use of Taylor series). For $\alpha,\beta>-1/2$ this error degradation is not of much concern because the extreme weights are the smallest; however, this is not the case if $\alpha$ and/or $\beta$ are smaller than 1/2-1/2, when the extreme nodes are the most significant. For this reason, the algorithm recalculates those extreme nodes and weights in the angular variable $\theta=\arccos x$ for negative values of $\alpha$ or $\beta$. We refer to [13, Section 4] for further details.
Figure 4: Accuracy of the algorithm for computing the weights of Gauss-Jacobi quadrature as functions of $n$ and $\sqrt{\alpha^{2}+\beta^{2}}$ for two differrent accuracy measures: maximum relative error (left) and relative total error (right). The different colors correspond to different error interval: $(0,10^{-14}]$ for black, $(10^{-14},10^{-13}]$ for green, $(10^{-13},10^{-12}]$ for blue, $(10^{-12},10^{-11}]$ for magenta dots, $(10^{-11},10^{-10}]$ for orange and $(10^{-10},10^{-9}]$ for cyan. At red dots the GW algorithm fails to compute all the weights. In the right graph, no orange, cyan or red dots appear.
@@
Our methods produce results with nearly double precision relative accuracy for all the nodes except possibly one. One has to take into accout that one node could be exactly zero or close to zero and that relative accuracy is not possible for that node. This situation occurs trivially for odd-degree Gauss-Gegenbauer quadrature but also for Jacobi non-symmetrical cases. This is an unavoidable limitation in the computation of the nodes. We notice, however, that we may have this limitation for one node at most, differently to the approach in [14], where all the nodes are computed with absolute accuracy.
@@
Figure 5: Accuracy of the algorithm for computing the nodes of Gauss-Jacobi quadrature for two differrent accuracy measures: maximum relative error (left) and second maximum relative error (right). The different colors correspond to different error intervals: $(0,10^{-15}]$ for black, $(10^{-15},10^{-14}]$ for green, $(10^{-14},10^{-13}]$ for blue, $10^{-13},10^{-12})$ for magenta, $10^{-12},10^{-11})$ for orange and $10^{-11},10^{-10})$ for cyan. Magenta, orange and cyan dots are represented with wider circles for visibility.
@@
The iterative method is faster that the asymptotic method by around a factor two, and this is why we use the iterative method when both give similar accuracy. On the other hand, our asymptotic method is faster than methods described in [14, 3] because our expansions (particularly for the nodes) are sufficiently accurate so the iterative refinement is not needed. As a result, our algorithm appears to be significantly faster than previous approaches, as Fig. 6 shows.
Figure 6: CPU times per node spent by the algorithms legpts (red dots) and legen (back dots) as a function of the degree $n$. The times are computed with Matlab version 2024 running in a laptop with Intel i7-1355U 1.70 GHz processor.
@@
3 Gauss-Laguerre quadrature
As for the Jacobi case, the algorithms for Gauss-Laguerre quadratures combine purely iterative and purely asymptotic methods, applying each of them in complementary regions. The iterative method is described in [10] and the asymptotic approximations are developed in [9]. In this section, we outline the main ingredients of the iterative method and we describe further details of the asymptotic method that were not fully developed in [9] and are necessary in the construction of the algorithms. Before this we recall some basic properties.
@@ 3 Gauss-Laguerre quadrature
where $z_{i}=\sqrt{x_{i}}$. We call the quantities $\omega_{i}$ scaled weights. The scaled weights have slow variation as a function of $i$ compared with the unscaled weights $w_{i}$. This is in accordance with the asymptotic estimate for the weights [21, 15.3.15]
@@ 3.1 Iterative method
For the Gauss-Laguerre case, the iterative method is the main procedure for computing the nodes and weights. We again employ the method for second order ODEs described in [20], using the ODE (14). All the computations necessary can be performed in the variable $z$ corresponding to the ODE (14), including local Taylor series for computing $y(z)$ and its derivative, and no other variables are required (differently from the Gauss-Jacobi case).
The ODE (14) is optimal for several reasons. In the first place, the coefficient $A(z)$ has simple monotonicity properties, which is important because the fixed point method sweep of [20] should be applied in the direction of decreasing $A(z)$. On the other hand, for large $n$ the coefficient $A(z)$ has small variation in most of the oscillation region, and because the nonlinear method of [20] is exact for ODEs with constant coefficients, the speed of convergence for computing each node tends to improve as $n$ increases. In the third place, the scaled weights $\omega_{i}$ are well conditioned as a function of $z_{i}$ because $\ddot{y}(z_{i})=0$. Finally, the resulting method admits subsampling in a natural way and one can easily compute only the weights that are greater of a given threshold.
@@
controls the size of the generated weights $w_{i}$ with respect to the dominant weight $w_{s}$. If we want to compute only those weights such that $w_{i}/w_{s}>\delta$ we can stop the algorithm when the index $i$ is such that $F_{i}<\log(\delta)$. Typically, there are two such values of $i$ for $\alpha$ large, one smaller and one large than $s$. Subsampling can safe a considerable amount of time, particularly for large degrees, because the weights decay exponentially, as (16) shows.
@@ 3.2.1 Airy-type expansion
The asymptotic computation of the nodes with the Airy expansion is described in detail in [11, Section 3.2.1]. The starting point is the following representation for $L_{n}^{(\alpha)}(\nu x)$ discussed in section 3.2 of [9] (see also [5])
@@ 3.2.2 Bessel-type expansion
The evaluation of the weights involves the asymptotic computation (for large $n$) of the function $\dot{y}(z_{i})$ in (15) by means of the expansion of [9, Eqs. (114) and (115)] [^4] We notice that we missed to mention in Eq. (115) of [9] that the variable $\varphi$ appearing in this equation is $\varphi=\frac{1}{2}(1-1/\rho)^{1/2}/\sqrt{\zeta}$., and this requires the computation of the Bessel function $J_{\alpha}\left(\nu\sqrt{\zeta}\right)$ for values of $\nu\sqrt{\zeta}$ near zeros of $J_{\alpha}\left(\nu\sqrt{\zeta}\right)$ [^5] Observe that for the Bessel-type expansions in [9, sections 3.3 and 3.4] we denoted $\nu=2n+\alpha+1$, while here we prefer to denote $\nu=4n+2\alpha+2$, as in Eq. (14), in order to be consistent with the notation of the Airy expansion we described earlier.. This computation can be unstable if done directly using standard numerical implementations for the Bessel function. The situation is very similar to the one we described before for the Airy-type expansion.
@@
and $\delta_{k}$ is accurately computed. Then, for computing $J_{\alpha}(j_{k}+\delta_{k})$, where $J_{\alpha}(j_{k})=0$ and $\delta_{k}/j_{k}<<1$ we proceed similarly as in [11, section 3.3], using the expansion

### v2

@@ 2.3 General Jacobi case
The methods considered for the general Gauss–Jacobi case combine the iterative method of [17] with the asymptotic methods of [15]. The iterative method provides a fast method of computation for a wide range of parameters; however, the accuracy of the weights worsens gradually as the degree increases and the asymptotic methods can be used as a replacement for moderate values of the parameters $\alpha$ and $\beta$.
@@ 2.3.1 Iterative method
The previous theorem gives a procedure to compute successive zeros in the direction of decreasing values of $\Omega(z)$. For a simple application of the method it is convenient that the monotonicity properties of $\Omega(z)$ are available and that the variation of the coefficient $\Omega(z)$ is as small as possible, because the method is exact if $\Omega(z)$ is constant. Different changes of variable can be considered such that the transformed differential equation becomes suitable in this sense. This has to be combined with an adequate method for computing the Jacobi polynomial and its derivative (or related functions [^3] Notice that the fixed point method applies to ODEs in normal form, and that transforming the ODE to this form may require a change of function, as for example in Eq. (1)); in our algorithms, as in [17], the main method considered for this purpose is Taylor series, with the successive derivatives in the series computed by differentiation of the ODE satisfied by Jacobi polynomials (or related functions). This approach was earlier considered in [18], and it is a simple but effective method of computation which, however, needs to be complemented with some alternative methods in the Jacobi case (and also for Gauss–Laguerre, as we later discuss).
The algorithm for the general Jacobi case combines the use of three different changes of variable. For computing the Jacobi polynomial and its derivative (or related functions) with Taylor series the original variable $x$ is used because the differential equation has polynomial coefficients and it is therefore easy to differentiate. The principal change of variable considered for the applying the fixed point method is $z=\tanh^{-1}(x)$; the coefficient of the Liouville-transformed ODE in normal form in the variable $z$ has even simpler monotonicity properties than that of Eq. (2). Finally, for nodes close to the end points $\pm 1$ the angular variable $\theta=\arccos x$ is used to refine the results obtained with the fixed point method associated to the principal change of variable.
For the particular case of Gauss–Gegenbauer quadrature we have implemented an arbitrary precision Maple algorithm in addition to the MATLAB algorithm. The computation of the Gegenbauer polynomial and its derivative (or related functions) by Taylor series is particularly simple due to the symmetry of the problem. The angular variable is not used in this Maple algorithm.
For further details on the iterative method for Gauss–Jacobi quadrature, we refer to [17].
@@ 2.3.2 Asymptotic method
Asymptotic methods have been employed for the computation of Gauss–Jacobi quadratures both in combination with iterative methods as in [20] or as independent iteration-free methods for Gauss–Legendre [1] and Gauss–Jacobi [15]. The Gauss–Hermite and Gauss–Laguerre cases were also described in [13]. In [26], a Riemann-Hilbert analysis is considered for the three classical cases and some extensions.
In our algorithms, the asymptotic computation of Gauss–Jacobi quadratures is based on the two types of asymptotic expansions described in [15]: an expansion in terms of elementary functions and a Bessel-type expansion [6]. The elementary expansion is valid in any fixed compact subinterval of (1,1)(-1,1) (therefore the nodes closest to $\pm 1$ are excluded). Contrarily, the Bessel expansion is uniformly valid in $1+\delta,1]$, where $\delta$ is an arbitrarily small positive number; because we can use the symmetry property $P_{n}^{(\alpha,\beta)}(x)=(-1)^{n}P_{n}^{(\beta,\alpha)}(-x)$ this expansion can be used for computing all the zeros of the Jacobi polynomials in (1,1)(-1,1).
The Bessel-type expansion is the main method of computation in our algorithms. However, the expansion in terms of elementary functions is convenient for computing the nodes with small absolute value in order to maintain relative accuracy, and not only absolute accuracy, similarly as discussed earlier for Gauss–Legendre quadrature. In our algorithm, $10\%$ of the nodes are computed with the elementary expansion and the rest with the Bessel-type expansion.
@@ 2.3.3 Combining the iterative and asymptotic methods
Our algorithm computes Gauss–Jacobi quadratures using the iterative method and the asymptotic expansions in complementary regions. The iterative method is fast and globally convergent and can be used for practically any values of the parameters, but with some error degradation as the degree increases. For that reason, it is convenient to replace the iterative method by asymptotic computation for sufficiently large degrees.
@@
The accuracy comparison between the asymptotic and iterative methods is shown in Fig. 2, where accuracy comparisons are performed for $10^{4}$ points randomly and uniformly generated in the intervals $n\in[1,3000]$, $\alpha,\beta\in(-1,15]$. The iterative method, as we will see later, turns out to be around a factor $2$ faster than the asymptotic method and, for this reason, when the accuracy of both methods is similar we prefer to use the iterative method. With this consideration, and given the accuracy checks shown in Fig. 2, a good compromise consist in using the asymptotic method when $n>250$ and $\alpha^{2}+\beta^{2}<4(n-195)/55$, and the iterative method otherwise. To this, we add the restriction $\alpha^{2}+\beta^{2}<2400$ in order to avoid overflow/underflow problems in the computation of the asymptotic expansions (not shown in Fig. 2).
Figure 2: Comparison of the accuracy of the iterative and the asymptotic method for Gauss–Jacobi quadratures, as a function of the degree $n$ and $\alpha^{2}+\beta^{2}$, for three different measures of accuracy: average relative error (left), maximum relative error (center) and relative total error (right). Denoting by $R$ the ratio of the accuracy of the asymptotic method over the accuracy of the iterative method, we have $R<1$ at the red points. $1\leq R<10$ at the green points and $R\geq 10$ at the blue points. The separation line $\alpha^{2}+\beta^{2}=\min\{\frac{\displaystyle{4}}{\displaystyle{55}}(n-195)\}$ is also shown.
@@ 2.3.4 Testing the combined algorithm
We now analyze the performance of the proposed algorithm in terms of computational range, accuracy, and speed. As we demonstrate below, it outperforms previous methods by being faster and by remaining reliable over a wider range of parameter values, including large $\alpha$ and $\beta$.
@@
As shown in Fig. 3, our algorithm is able to compute Gauss–Jacobi quadratures parameters for $\alpha,\,\beta\in(-1,1000]$ and $n\leq 10^{6}$, while for larger values of $\alpha$ and $\beta$ the algorithm may fail to produce results. For generating the points for the figure we have considered values of $n$ with $n=\lfloor 10^{6R}\rfloor$ with $R$ uniform pseudorandom numbers in $[0,1]$; $\alpha$ and $\beta$ are also generated by exponentiation of pseudorandom numbers.
@@
The maximum relative error when the iterative method is applied usually takes place at the extremes of the interval, which corresponds with the less significant weights when $\alpha,\beta>-1/2$ (see Eq. (4)). This is expected because these are the last computed weights in the iterative algorithm and they are the closest to the singularities of the ODE (which affects the use of Taylor series). For $\alpha,\beta>-1/2$ this error degradation is not of much concern because the extreme weights are the smallest; however, this is not the case if $\alpha$ and/or $\beta$ are smaller than 1/2-1/2, when the extreme nodes are the most significant. For this reason, the algorithm recalculates those extreme nodes and weights in the angular variable $\theta=\arccos x$ for negative values of $\alpha$ or $\beta$. We refer to [17, Section 4] for further details.
Figure 4: Accuracy of the algorithm for computing the weights of Gauss–Jacobi quadrature as functions of $n$ and $\sqrt{\alpha^{2}+\beta^{2}}$ for two different accuracy measures: maximum relative error (left) and relative total error (right). The different colors correspond to different error interval: $(0,10^{-14}]$ for black, $(10^{-14},10^{-13}]$ for green, $(10^{-13},10^{-12}]$ for blue, $(10^{-12},10^{-11}]$ for magenta dots, $(10^{-11},10^{-10}]$ for orange and $(10^{-10},10^{-9}]$ for cyan. At red dots the GW algorithm fails to compute all the weights. In the right graph, no orange, cyan or red dots appear.
@@
Our methods produce results with nearly double precision relative accuracy for all the nodes except possibly one. One has to take into account that one node could be exactly zero or close to zero and that relative accuracy is not possible for that node. This situation occurs trivially for odd-degree Gauss–Gegenbauer quadrature but also for Jacobi non-symmetrical cases. This is an unavoidable limitation in the computation of the nodes. We notice, however, that we may have this limitation for one node at most, differently to the approach in [20], where all the nodes are computed with absolute accuracy.
@@
Figure 5: Accuracy of the algorithm for computing the nodes of Gauss–Jacobi quadrature for two differrent accuracy measures: maximum relative error (left) and second maximum relative error (right). The different colors correspond to different error intervals: $(0,10^{-15}]$ for black, $(10^{-15},10^{-14}]$ for green, $(10^{-14},10^{-13}]$ for blue, $10^{-13},10^{-12})$ for magenta, $10^{-12},10^{-11})$ for orange and $10^{-11},10^{-10})$ for cyan. Magenta, orange and cyan dots are represented with wider circles for visibility.
@@
The iterative method is faster than the asymptotic method by around a factor two, and this is why we use the iterative method when both give similar accuracy. On the other hand, our asymptotic method is faster than methods described in [20, 4] because our expansions (particularly for the nodes) are sufficiently accurate so the iterative refinement is not needed. As a result, our algorithm appears to be significantly faster than previous approaches, as Fig. 6 shows.
Figure 6: CPU times in seconds per node and weight spent by the algorithms jacpts (red dots) and GJ (black dots) as a function of the degree $n$. The times are computed with MATLAB version 2024b running in a laptop with Intel(R) Core(TM) i7-1355U processor.
@@
3 Gauss–Laguerre quadrature
As for the Jacobi case, the algorithms for Gauss–Laguerre quadratures combine purely iterative and purely asymptotic methods, applying each of them in complementary regions. The iterative method is described in [14] and the asymptotic approximations are developed in [13]. In this section, we outline the main ingredients of the iterative method and we describe further details of the asymptotic method that were not fully developed in [13] and are necessary in the construction of the algorithms. Before this we recall some basic properties.
@@ 3 Gauss–Laguerre quadrature
where $z_{i}=\sqrt{x_{i}}$. We call the quantities $\omega_{i}$ scaled weights. The scaled weights have slow variation as a function of $i$ compared with the unscaled weights $w_{i}$. This is in accordance with the asymptotic estimate for the weights [29, 15.3.15]
@@ 3.1 Iterative method
For the Gauss–Laguerre case, the iterative method is the main procedure for computing the nodes and weights. We again employ the method for second order ODEs described in [27], using the ODE (15). All the computations necessary can be performed in the variable $z$ corresponding to the ODE (15), including local Taylor series for computing $y(z)$ and its derivative, and no other variables are required (differently from the Gauss–Jacobi case).
The ODE (15) is optimal for several reasons. In the first place, the coefficient $A(z)$ has simple monotonicity properties, which is important because the fixed point method sweep of [27] should be applied in the direction of decreasing $A(z)$. On the other hand, for large $n$ the coefficient $A(z)$ has small variation in most of the oscillation region, and because the nonlinear method of [27] is exact for ODEs with constant coefficients, the speed of convergence for computing each node tends to improve as $n$ increases. In the third place, the scaled weights $\omega_{i}$ are well conditioned as a function of $z_{i}$ because $\ddot{y}(z_{i})=0$. Finally, the resulting method admits subsampling in a natural way and one can easily compute only the weights that are greater than a given threshold.
@@
controls the size of the generated weights $w_{i}$ with respect to the dominant weight $w_{s}$. If we want to compute only those weights such that $w_{i}/w_{s}>\delta$ we can stop the algorithm when the index $i$ is such that $F_{i}<\log(\delta)$. Typically, there are two such values of $i$ for $\alpha$ large, one smaller and one large than $s$. Subsampling can save a considerable amount of time, particularly for large degrees, because the weights decay exponentially, as (17) shows.
@@ 3.2.1 Airy-type expansion
The asymptotic computation of the nodes with the Airy expansion is described in detail in [15, Section 3.2.1]. The starting point is the following representation for $L_{n}^{(\alpha)}(\nu x)$ discussed in section 3.2 of [13] (see also [7])
@@ 3.2.2 Bessel-type expansion
The evaluation of the weights involves the asymptotic computation (for large $n$) of the function $\dot{y}(z_{i})$ in (16) by means of the expansion of [13, Eqs. (114) and (115)] [^5] We notice that we missed to mention in Eq. (115) of [13] that the variable $\varphi$ appearing in this equation is $\varphi=\frac{1}{2}(1-1/\rho)^{1/2}/\sqrt{\zeta}$., and this requires the computation of the Bessel function $J_{\alpha}\left(\nu\sqrt{\zeta}\right)$ for values of $\nu\sqrt{\zeta}$ near zeros of $J_{\alpha}\left(\nu\sqrt{\zeta}\right)$ [^6] Observe that for the Bessel-type expansions in [13, sections 3.3 and 3.4] we denoted $\nu=2n+\alpha+1$, while here we prefer to denote $\nu=4n+2\alpha+2$, as in Eq. (15), in order to be consistent with the notation of the Airy expansion we described earlier.. This computation can be unstable if done directly using standard numerical implementations for the Bessel function. The situation is very similar to the one we described before for the Airy-type expansion.
@@
and $\delta_{k}$ is accurately computed. Then, for computing $J_{\alpha}(j_{k}+\delta_{k})$, where $J_{\alpha}(j_{k})=0$ and $\delta_{k}/j_{k}<<1$ we proceed similarly as in [15, section 3.3], using the expansion